When x2 - 3x + 2k is divided by x +4, the remainder is 21. Find k k= (Simplify your answer.)
Answer #1

2 giveni Let f(x)= x -3x+qk is divided by (x+4) remainder is 21 then at x=-4 f(-4) = 21 +4)2-3(-4)+2K =2 16+ 12+2k=21 ak= 21-